# Periodic orbits and their gravitational wave radiations in black hole with dark matter halo

### v1

@@ Abstract
This study investigates the gravitational dynamics of a spherically symmetric black hole embedded within a dark matter halo, focusing on the impact of the dark matter halo on time-like particle geodesics. We analyze key orbital parameters, including marginally bound orbits, innermost stable circular orbits, and periodic orbits, demonstrating how the dark matter halo alters their energy and angular momentum compared to an isolated black hole. Furthermore, we calculate the gravitational waveforms emitted by these periodic orbits, providing potential observational signatures for such systems. Our results demonstrate a clear correlation between the gravitational waveforms emitted by a small object orbiting a supermassive black hole and the object’s zoom-whirl orbital behavior. Higher zoom numbers correspond to more complex waveform substructures. the presence of a dark matter halo significantly impacts these waveforms. This research contributes to a deeper understanding of the interplay between black holes and dark matter halos, offering insights into their complex gravitational interactions.
@@ I Introduction
The landmark detection of gravitational waves (GWs) from binary black hole mergers by LIGO and Virgo LIGOScientific:2016aoc ; LIGOScientific:2016vbw ; LIGOScientific:2016vlm ; LIGOScientific:2016emj has opened a new era for probing strong-field gravity and black hole dynamics. These detections have not only validated predictions of general relativity (GR) in highly nonlinear regimes but have also opened new avenues for studying compact objects and their astrophysical environments. Among the most promising targets for future space-based GW detectors are extreme mass ratio inspirals (EMRIs), where a stellar-mass compact object, such as a black hole or neutron star, gradually spirals into a supermassive black hole (SMBH). Space-based GW detectors like LISA LISA:2022yao ; LISA:2017pwj , Taiji Hu:2017mde ; Gong:2021gvw , TianQin Liu:2020eko ; TianQin:2015yph ; Gong:2021gvw , and DECIGO Musha:2017usi are ideally suited to detect the long-duration, low-frequency signals emitted by these systems. EMRIs are of particular significance because their gravitational waveforms encode detailed information about the spacetime structure of the central black hole Hughes:2000ssa ; Glampedakis:2005hs ; Barausse:2020rsu . This provides an unparalleled opportunity to probe the Kerr geometry predicted by GR and to search for potential deviations caused by new physics, such as modified gravity or the presence of exotic compact objects Cardoso:2019rvt ; Babak:2017tow .
@@
A critical feature of EMRI dynamics is the prevalence of periodic orbits--bound trajectories in which a particle returns to its initial state after an integer number of radial and angular oscillations Levin:2008ci ; Levin:2008mq . These orbits, indexed by zoom--whirl parameters $(z,w,v)$ Levin:2008ci ; Levin:2008mq ; Levin:2009sk , serve as foundational modes for understanding generic orbital dynamics Levin:2008mq ; Levin:2009sk ; Misra:2010pu ; Babar:2017gsg . The framework of periodic orbits has been extensively applied to a variety of black hole spacetimes, including charged black holes Misra:2010pu , naked singularities Babar:2017gsg , Kerr-Sen black holes Liu:2018vea , and Einstein-Lovelock black holes Lin:2021noq , Kehagias-Sfetsos black holes Wei:2019zdf , quantum-corrected black holes Deng:2020yfm and brane-world black holes Deng:2020hxw . For further studies of periodic orbits in other black hole spacetimes, see Refs. Lin:2023eyd ; Chan:2025ocy ; Zhou:2020zys ; Lin:2022wda ; Zhang:2022psr ; Lin:2022llz ; Habibina:2022ztd ; Zhang:2022zox ; Wang:2022tfo ; Lin:2023rmo ; Yao:2023ziq ; Tu:2023xab ; Yang:2024lmj ; Shabbir:2025kqh ; Junior:2024tmi ; Zhao:2024exh ; Jiang:2024cpe ; Yang:2024cnd ; Meng:2024cnq ; Li:2024tld ; QiQi:2024dwc . Gravitational waveforms from EMRIs inherit distinct signatures of these periodic orbits, particularly the characteristic “zoom-whirl” phases during the inspiral Tu:2023xab ; Yang:2024lmj . This has motivated numerous studies on the GWs emitted by periodic orbits in various black hole spacetimes, as well as their potential detectability by future space-based GW observatories Tu:2023xab ; Yang:2024lmj ; Shabbir:2025kqh ; Junior:2024tmi ; Zhao:2024exh ; Jiang:2024cpe ; Yang:2024cnd ; Meng:2024cnq ; Li:2024tld ; QiQi:2024dwc . However, such dynamics are not isolated from the galactic environment: dark matter (DM) halos, which dominate the mass distribution in galaxies Navarro:1995iw ; Hernquist:1990be ; Gondolo:1999ef , can imprint measurable modifications on the SMBH’s spacetime Cardoso:2021wlq ; Sadeghian:2013laa ; Barausse:2014tra .
@@
DM halos, integral components of galaxies, significantly influence the dynamics of objects within them. These halos are often modeled using various density profiles, each designed to capture the diverse characteristics of galaxies. A generalized parametrization, frequently employed, expresses the density $\rho(r)$ as a function of radial distance ($r$) from the halo’s center Taylor:2002zd
@@
where $a_{0}$ represents the length-scale of the galaxy, $\rho_{0}$ is the DM density at $r=a_{0}$, and $\alpha$, $\gamma$, and $k$ are parameters defining the specific profile. These parameters reflect variations in the galaxy’s size, mass, and shape. The profile plays a crucial role in shaping the gravitational potential near SMBHs Baes:2004pg .
[…]
It is natural to explore that how the DM distributions can affect the spacetime of the central black hole. One approach involves modeling an isolated black hole spacetime matched to a specific matter distribution via the mass function Hou:2018bar ; Xu:2018wow ; Konoplya:2019sns ; Jusufi:2020cpn ; Liu:2021xfb ; Zhang:2022roh ; Zhang:2021bdr ; Xu:2021dkv . Another method employs relativistic extensions of Einstein clusters Einstein_1939 ; Cardoso:2021wlq ; Ciotti:1996pf , which provide a framework for constructing self-consistent spacetimes of SMBHs embedded in DM halos. In such models, the anisotropic pressure of the halo modifies the effective potential, affecting key orbital properties such as the innermost stable circular orbits (ISCOs) and light rings. For instance, Cardoso et al. Cardoso:2021wlq demonstrated that DM halos can induce gravitational redshift and alter GW propagation, resulting in detectable phase shifts in EMRIs’ signals. In their study, Cardoso et al. Cardoso:2021wlq utilized a generalized version of the ”Einstein cluster” to develop an exact field solution for a black hole immersed in DM. Their model adopted a Hernquist-type density distribution for the DM halo, given by:
@@
Using similar methods, spacetime solutions for black holes embedded in DM with alternative density profiles have been constructed in a series of works Konoplya:2022hbl ; Jusufi:2022jxu ; Shen:2023erj ; Shen:2024qxv ; Maeda:2024tsg ; Al-Badawi:2024asn , extending the applicability of such models to a broader range of galactic environments.
While periodic orbits in vacuum black hole spacetimes are well-studied, the role of DM halos remains unexplored. In this paper, we analyze periodic orbits and their GW signatures for SMBHs surrounded by Hernquist-like DM halos Hernquist:1990be . Using the balck hole solution sorrounding by the DM halo derived in Cardoso:2021wlq from a generalized Einstein cluster model Einstein_1939 , we derive the effective potential for timelike geodesics and compute corrections to marginally bound orbits (MBOs) and ISCOs. We show that the halo’s length scale $a_{0}$ and central density $\rho_{0}$ reduce orbital radii and angular momenta, shifting the $(z,w,v)$ taxonomy of periodic orbits. Employing the numerical “kludge” framework Babak:2006uv , we simulate GW waveforms for representative orbits, revealing DM-induced phase shifts and amplitude modulations.
@@ II Black hole spacetime in DM halo
This section presents a brief introduction to a spherically symmetric black hole in the Hernquist-like DM halo. The line element of this black hole is given by Cardoso:2021wlq
@@
which, at small distances, describes a black hole with mass $M_{B}$, while at large scales, it corresponds to a distribution of matter described by Eq. (1.2). This spacetime, similar to the Schwarzschild black hole, possesses an event horizon at $r=2M_{BH}$ and a curvature singularity at $r=0$. This profile corresponds to matter density
@@
which indicates that for a fixed value of $M_{H}$ and large $a_{0}$ or for increasing $r$ the DM density $\rho_{\rm DM}$ decreases. The ADM mass of this black hole here is the sum of the mass of the DM halo and that of the black hole, i.e., $M_{ADM}=M_{H}+M_{B}$. It is also worth mentioning that the matter density vanishes at the horizon, and the tangential pressure is regular there. The ratio $M/a_{0}$ is termed the compactness parameter and can assume values greater than $10^{-4}$ in a galactic context Cardoso:2021wlq .
@@ III.2 Effective potential and bounded orbtis
is the radial effective potential that describes the radial motion of the particle. Increasing angular momentum changes the effective potential, creating extrema. This behavior is illustrated in Fig. 1 where it is observed that as the angular momentum increases, the value of $V_{\rm eff}$ also increases. For fixed $a_{0}$, as extrema of $V_{\rm eff}$ decreases with increasing $M_{H}$. Conversely, for a fixed $M_{H}$, the extrema of $V_{\rm eff}$ increase with increasing $a_{0}$. Since the function $a(r)$ admits asymptotic flatness, thus $V_{\rm eff}\rightarrow 1$ as $r\rightarrow+\infty$. So $E=1$ is the critical energy of the particle moving in an orbit. The particle having energy $E>1$ will move with positive velocity at infinity i.e. $\dot{r}^{2}>0$ as $E^{2}-{V_{\rm eff}}_{r\rightarrow+\infty}>0$, implying the existence of unbounded orbits. Therefore, the maximum energy the particle can acquire while moving in a bounded orbit must be $E\leq 1$.
@@
For a massive particle around a spherically symmetric black hole immersed in DM halo, bounded orbits exist between MBOs and ISCOs. Between these two points, there exists a range of energies that allow for bound orbits to exist. The energy of a particle in a bound orbit around a black hole must indeed lie within the range $E_{\rm ISCO}\leq E\leq E_{\rm MBO}=1$. $E_{\rm MBO}=1$ corresponds to the energy of the MBO, where the particle has just enough energy to escape to infinity. $E_{\rm ISCO}$ represents the energy of the ISCO, and any particle with energy less than this will plunge into the black hole. The condition for angular momentum is slightly different. The angular momentum $L$ of a particle in a bound orbit must satisfy $L\geq L_{\rm ISCO}$, where $L_{\rm ISCO}$ is the angular momentum of the particle in ISCO.
@@ III.3 Marginally stable orbit
Solving these equations yields the radius and angular momentum of the MBO. While exact solutions cannot be computed due to the complexity of the equations, approximated solutions are calculated by considering $M/a_{0}$ as a small quantity. We get the radius for the MBO $r_{\rm MBO}$ as
@@ III.4 Innermost stable circular orbits
where $\tilde{r}_{\text{\tiny ISCO}}=6M_{B}$ and $\tilde{L}_{\text{\tiny ISCO}}=2\sqrt{3}M_{B}$ are the particle’s radius and angular momentum at ISCO’s around the Schwarzschild black hole. In Fig. (3), for a fixed value of $M_{H}$, as $a_{0}$ increases, both $r_{\rm ISCO}$ and $E_{\rm ISCO}$ increase while $L_{\rm ISCO}$ decreases. Conversely, when $a_{0}$ is fixed, increasing $M_{H}$ leads to a decrease in $r_{\rm ISCO}$ and $E_{\rm ISCO}$, while an increase in $L_{\rm ISCO}$. The allowed $E$-$L$ regions for the existence of bound orbits are plotted in Fig. (4). For fixed values of halo’s mass $M_{H}$ and length-scale $a_{0}$, energy $E$ changes with an increase in angular momentum $L$. The width of these $E$-$L$ regions expands with increasing $M_{H}$ and contracts with increasing $a_{0}$.
@@
Figure 4: The allowed $L-E$ parametric space for a timelike particle orbiting around a black hole with DM halo. Upper panel: The value of $a_{0}$ is kept fixed at $a_{0}=1000M_{B}$ while ${M_{H}}/{M_{B}}$= 5,15, 25, 35, 45 (top to bottom). Lower panel: The value of $M_{H}=10M_{B}$ is kept fixed while $a_{0}/M_{B}$= 100, 200, 500, 1000, 5000 (bottom to top).
@@ IV Periodic Orbits
Figure 7: Periodic orbits for different values of $(z,w,v)$ around a black hole with DM halo. The value of parameters are $M_{H}=10M_{H}$, $a_{0}=1000M_{H}$ and $E=0.96$.
@@
Building on the taxonomy established in Levin:2008ci , we define the frequency ratio $q$ as the relationship between the radial ($\omega_{r}$) and azimuthal ($\omega_{\phi}$) oscillation frequencies. This ratio is expressed in terms of three integers $(z,w,v)$, zoom $z$, whirl $w$ and vertex $v$ as
@@
where $\Delta\phi$ is given by Eq. (4.1). For irrational values of $q$, the particle’s trajectory traces a precessing orbit, characterized by a precession angle $w=\Delta\Phi-2\pi$. Rational values of $q$, however, correspond to periodic orbits, where the particle returns to its initial position after a finite time. As shown in Levin:2008ci , generic orbits can be interpreted as perturbations of these periodic orbits. Thus, the study of periodic orbits provides critical insights into the behavior of generic orbits and the gravitational radiation emitted near black holes.
@@
Fig. 5 and Fig. 6 illustrate the behavior of the rational number $q$ with respect to energy $E$ and angular momentum $L$, respectively, for various black hole masses $M_{H}$. Fig. 5 reveals a gradual increase in $q$ with increasing $E$, while it reaches a limited value as energy reaches its maximum. This maximum energy exhibits an inverse relationship with $M_{H}$, decreasing as $M_{H}$ increases. Conversely, Fig. 6 demonstrates a gradual decrease in $q$ with increasing angular momentum $L$, ceasing at the minimum value of $L$. Furthermore, both figures indicate a direct relationship between $L$ and $E$ for a fixed $q$: as $E$ increases, $L$ also increases.
By fixing energy to $E=0.96$ in Fig. 7 and parameter $\epsilon=0.3$ in Fig. 8, periodic orbits are illustrated using different combinations of integers $(z,w,v)$. The value of $z$ determines the number of blades in the orbit’s shape. Larger $z$ values correspond to larger blade profiles and increasingly complex trajectories.

### v2

@@ Abstract
This study investigates the gravitational dynamics of a spherically symmetric black hole embedded within a dark matter halo, focusing on the impact of the dark matter halo on time-like particle geodesics. We analyze key orbital parameters, including marginally bound orbits, innermost stable circular orbits, and periodic orbits, demonstrating how the dark matter halo alters their energy and angular momentum compared to an isolated black hole. Furthermore, we calculate the gravitational waveforms emitted by these periodic orbits, providing potential observational signatures for such systems. Our results demonstrate a clear correlation between the gravitational waveforms emitted by a small object orbiting a supermassive black hole and the object’s zoom-whirl orbital behavior. Higher zoom numbers correspond to more complex waveform substructures. The presence of a dark matter halo significantly impacts these waveforms. We have conducted a detailed analysis of the frequency spectra of the gravitational waveforms arising from these periodic orbits and evaluated their potential detectability. This research contributes to a deeper understanding of the interplay between black holes and dark matter halos, offering insights into their complex gravitational interactions.
@@ I Introduction
The landmark detection of gravitational waves (GWs) from binary black hole mergers by LIGO and Virgo [1, 2, 3, 4] has opened a new era for probing strong-field gravity and black hole dynamics. These detections have not only validated predictions of general relativity (GR) in highly nonlinear regimes but have also opened new avenues for studying compact objects and their astrophysical environments. Among the most promising targets for future space-based GW detectors are extreme mass ratio inspirals (EMRIs), where a stellar-mass compact object, such as a black hole or neutron star, gradually spirals into a supermassive black hole (SMBH). Space-based GW detectors like LISA [5, 6], Taiji [7, 8], TianQin [9, 10, 8], and DECIGO [11] are ideally suited to detect the long-duration, low-frequency signals emitted by these systems. EMRIs are of particular significance because their gravitational waveforms encode detailed information about the spacetime structure of the central black hole [12, 13, 14]. This provides an unparalleled opportunity to probe the Kerr geometry predicted by GR and to search for potential deviations caused by new physics, such as modified gravity or the presence of exotic compact objects [15, 16].
@@
A critical feature of EMRI dynamics is the prevalence of periodic orbits-bound trajectories in which a particle returns to its initial state after an integer number of radial and angular oscillations [17, 18]. These orbits, indexed by zoom-whirl parameters $(z,w,v)$ [17, 18, 19], serve as foundational modes for understanding generic orbital dynamics [18, 19, 20, 21]. The framework of periodic orbits has been extensively applied to a variety of black hole spacetimes, including charged black holes [20], naked singularities [21], Kerr-Sen black holes [22], and Einstein-Lovelock black holes [23], Kehagias-Sfetsos black holes [24], quantum-corrected black holes [25] and brane-world black holes [26]. For further studies of periodic orbits in other black hole spacetimes, see Refs. [27, 28, 29, 30, 31, 32, 33, 34, 35, 37, 36, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. Gravitational waveforms from EMRIs inherit distinct signatures of these periodic orbits, particularly the characteristic “zoom-whirl” phases during the inspiral [38, 39]. This has motivated numerous studies on the GWs emitted by periodic orbits in various black hole spacetimes, as well as their potential detectability by future space-based GW observatories [38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. However, such dynamics are not isolated from the galactic environment: dark matter (DM) halos, which dominate the mass distribution in galaxies [49, 50, 51], can imprint measurable modifications on the SMBH’s spacetime [52, 53, 54].
@@
DM halos, integral components of galaxies, significantly influence the dynamics of objects within them. These halos are often modeled using various density profiles, each designed to capture the diverse characteristics of galaxies. A generalized parametrization, frequently employed, expresses the density $\rho(r)$ as a function of radial distance ($r$) from the halo’s center [55]
@@
where $a_{0}$ represents the length-scale of the galaxy, $\rho_{0}$ is the DM density at $r=a_{0}$, and $\alpha$, $\gamma$, and $k$ are parameters defining the specific profile. These parameters reflect variations in the galaxy’s size, mass, and shape. The profile plays a crucial role in shaping the gravitational potential near SMBHs [56].
Several well-established models utilize this generalized form. For example, the Navarro-Frenk-White (NFW) profile [49] commonly used for galaxies with significant DM content, sets $\alpha=1$, $\gamma=3$, and $k=1$. The NFW profile describes a cuspy density distribution near the center, gradually declining at larger radii. For dwarf galaxies, the Burkert profile, which corresponds to $\alpha=1$, $\gamma=3$, and $k=2$, is often more suitable [57]. This profile features a core-like structure in the central region, avoiding the central cusp of the NFW profile. Another common model is the Hernquist profile, which corresponds to $\alpha=1$, $\gamma=4$, and $k=1$, often used to describe the Seersic profiles observed in elliptical galaxies and galactic bulges [50].
It is natural to explore that how the DM distributions can affect the spacetime of the central black hole. One approach involves modeling an isolated black hole spacetime matched to a specific matter distribution via the mass function [58, 59, 60, 61, 62, 63, 64, 65]. Another method employs relativistic extensions of Einstein clusters [66, 52, 67], which provide a framework for constructing self-consistent spacetimes of SMBHs embedded in DM halos. In such models, the anisotropic pressure of the halo modifies the effective potential, affecting key orbital properties such as the innermost stable circular orbits (ISCOs) and light rings. For instance, Cardoso et al. [52] demonstrated that DM halos can induce gravitational redshift and alter GW propagation, resulting in detectable phase shifts in EMRIs’ signals. In their study, Cardoso et al. [52] utilized a generalized version of the “Einstein cluster” to develop an exact field solution for a black hole immersed in DM. Their model adopted a Hernquist-type density distribution for the DM halo, given by:
@@
Using similar methods, spacetime solutions for black holes embedded in DM with alternative density profiles have been constructed in a series of works [68, 69, 70, 71, 72, 73], extending the applicability of such models to a broader range of galactic environments. Numerical approach to black hole solutions with generic dark matter profiles and their implications in GW radiations have also been explored in ref. [74].
While periodic orbits in vacuum black hole spacetimes are well-studied, the role of DM halos remains unexplored. In this paper, we analyze periodic orbits and their GW signatures for SMBHs surrounded by Hernquist-like DM halos [50]. Using the balck hole solution sorrounding by the DM halo derived in [52] from a generalized Einstein cluster model [66], we derive the effective potential for timelike geodesics and compute corrections to marginally bound orbits (MBOs) and ISCOs. We show that the halo’s length scale $a_{0}$ and central density $\rho_{0}$ reduce orbital radii and angular momenta, shifting the $(z,w,v)$ taxonomy of periodic orbits. Employing the numerical “kludge” framework [75], we simulate GW waveforms for representative orbits, revealing DM-induced phase shifts and amplitude modulations. We then conduct a detailed analysis of the frequency spectra of the gravitational waveforms arising from these periodic orbits and evaluated their potential detectability by comparing their characteristic strains with the sensitivity curve of space-based GW detector, LISA.
@@ II Black hole spacetime in DM halo
This section presents a brief introduction to a spherically symmetric black hole in the Hernquist-like DM halo. The line element of this black hole is given by [52, 74]
@@
which, at small distances, describes a black hole with mass $M_{B}$ while at large scales it corresponds to a distribution of matter described by Eq. (1.2). This spacetime, similar to the Schwarzschild black hole, possesses an event horizon at $r=2M_{BH}$ and a curvature singularity at $r=0$. This profile corresponds to matter density
@@
which indicates that for a fixed value of $M_{H}$ and large $a_{0}$ or for increasing $r$ the DM density $\rho_{\rm DM}$ decreases. The ADM mass of this black hole here is the sum of the mass of the DM halo and that of the black hole, i.e. $M_{ADM}=M_{H}+M_{B}$. It is also worth mentioning here that the matter density vanishes at the horizon and the tangential pressure is regular there. The ratio $M/a_{0}$ is termed the compactness parameter and can assume values greater than $10^{-4}$, in a galactic context [52].
@@ III.2 Effective potential and bounded orbtis
is the radial effective potential which describes the radial motion of the particle. Increasing angular momentum changes the effective potential, creating extrema. This behavior is illustrated in Fig. 1 where it is observed that as the angular momentum increases, the value of $V_{\rm eff}$ also increases. For fixed $a_{0}$, as extrema of $V_{\rm eff}$ decreases with increasing $M_{H}$. Conversely, for a fixed $M_{H}$, the extrema of $V_{\rm eff}$ increase with increasing $a_{0}$. Since the function $a(r)$ admits asymptotic flatness, thus $V_{\rm eff}\rightarrow 1$ as $r\rightarrow+\infty$. So $E=1$ is the critical energy of the particle moving in an orbit. The particle having energy $E>1$ will move with positive velocity at infinity i.e. $\dot{r}^{2}>0$ as $E^{2}-{V_{\rm eff}}_{r\rightarrow+\infty}>0$, implying the existence of unbounded orbits. Therefore the maximum energy the particle can acquire while moving in a bounded orbit must be $E\leq 1$.
@@
For a massive particle around a spherically symmetric black hole immersed in DM halo, bounded orbits exist between MBOs and ISCOs. Between these two points, there exists a range of energies that allow for bound orbits to exist. The energy of a particle in a bound orbit around a black hole must indeed lie within the range $E_{\rm ISCO}\leq E\leq E_{\rm MBO}=1$. $E_{\rm MBO}=1$ corresponds to the energy of the MBO, where the particle has just enough energy to escape to infinity. $E_{\rm ISCO}$ represents the energy of the ISCO, and any particle with energy less than this will plunge into the black hole. The condition for angular momentum is slightly different. The angular momentum $L$ of a particle in a bound orbit must satisfy $L\geq L_{\rm ISCO}$, where $L_{\rm ISCO}$ being the angular momentum of the particle in ISCO.
@@ III.3 Marginally stable orbit
Solving these equations yields the radius and angular momentum of the MBO. While exact solutions cannot be computed due to the complexity of the equations, thus approximated solutions are calculated by considering $M/a_{0}$ as a small quantity, we get the radius for the MBO $r_{\rm MBO}$ as
@@ III.4 Innermost stable circular orbits
where $\tilde{r}_{\text{\tiny ISCO}}=6M_{B}$ and $\tilde{L}_{\text{\tiny ISCO}}=2\sqrt{3}M_{B}$ are the particle’s radius and angular momentum at ISCO’s around the Schwarzschild black hole. In Fig. 3, for a fixed value of $M_{H}$, as $a_{0}$ increases, both $r_{\rm ISCO}$ and $E_{\rm ISCO}$ increase while $L_{\rm ISCO}$ decreases. Conversely, when $a_{0}$ is fixed, increasing $M_{H}$ leads to a decrease in $r_{\rm ISCO}$ and $E_{\rm ISCO}$, while an increase in $L_{\rm ISCO}$. The allowed $E$-$L$ regions for the existence of bound orbits are plotted in Fig. 4. For fixed values of halo’s mass $M_{H}$ and length-scale $a_{0}$ energy $E$ changes with increase in angular momentum $L$. The width of these $E$-$L$ regions expands with increasing $M_{H}$ and contracts with increasing $a_{0}$.
@@
Figure 4: The allowed LEL-E parametric space for a timelike particle orbiting around a black hole with DM halo. Upper panel: The value of $a_{0}$ is kept fixed at $a_{0}=1000M_{B}$ while ${M_{H}}/{M_{B}}$= 5,15, 25, 35, 45 (top to bottom). Lower panel: The value of $M_{H}=10M_{B}$ is kept fixed while $a_{0}/M_{B}$= 100, 200, 500, 1000, 5000 (bottom to top).
@@ IV Periodic Orbits
Figure 7: Periodic orbits for different values of $(z,w,v)$ around a black hole with DM halo. The value of parameters are $M=10M_{H}$, $a_{0}=1000M_{H}$ and $E=0.96$.
@@
Building on the taxonomy established in [17], we define the frequency ratio $q$ as the relationship between the radial ($\omega_{r}$) and azimuthal ($\omega_{\phi}$) oscillation frequencies. This ratio is expressed in terms of three integers $(z,w,v)$, zoom $z$, whirl $w$ and vertex $v$ as
@@
where $\Delta\phi$ is given by Eq. (4.1). For irrational values of $q$, the particle’s trajectory traces a precessing orbit, characterized by a precession angle $w=\Delta\Phi-2\pi$. Rational values of $q$, however, correspond to periodic orbits, where the particle returns to its initial position after a finite time. As shown in [17], generic orbits can be interpreted as perturbations of these periodic orbits. Thus, the study of periodic orbits provides critical insights into the behavior of generic orbits and the gravitational radiation emitted near black holes.
@@
Fig. 5 and Fig. 6 illustrate the behavior of the rational number $q$ with respect to energy $E$ and angular momentum $L$, respectively, for various black hole masses $M_{H}$. Fig. 5 reveals a gradual increase in $q$ with increasing $E$, while it reaches a limited value as energy reaches its maximum. This maximum energy exhibits an inverse relationship with $M_{H}$, decreasing as $M_{H}$ increases. Conversely, Fig. 6 demonstrates a gradual decrease in $q$ with increasing angular momentum $L$, ceasing at the minimum value of $L$. Furthermore, both figures indicate a direct relationship between $L$ and $E$ for a fixed $q$: as $E$ increases, $L$ also increases. In general, the range of $q$ depends on the value of angular momentum $L$. As we mentioned above, when one fixed $L$, $q$ reaches its maximum value as energy reaches its maximum [18]. It is interesting to note that when $L=L_{\rm MBO}$, the orbit with maximum energy corresponds to the homoclinic orbit [18], which has divergent number of whirls, indicating $q\to+\infty$.
To illustrate the periodic orbits and their GW radiations, we focus on certain periodic orbits with relatively small values of $q$ for two main reasons. First, high $q$ orbits correspond to extreme “whirl-dominated” motion near ISCO. These orbits in general are short-lived due to rapid gravitational wave radiation, thus are not expected to be astrophysical relevant. Second, orbits with large $q$ require finely tuned initial conditions and are highly sensitive to perturbations, such as those caused by gravitational wave back-reaction. This sensitivity is evident from Fig. 5 and Fig. 6, where small changes in energy or angular momentum lead to rapid increases in $q$. Then by fixing energy to $E=0.96$ in Fig. 7 and parameter $\epsilon=0.3$ in Fig. 8, periodic orbits are illustrated using different combinations of integers $(z,w,v)$. The value of $z$ determines the number of blades in the orbit’s shape. Larger $z$ values correspond to larger blade profiles and increasingly complex trajectories.